P(8,5)
_____________
P(3,1)x P(5,4)

Answer by checkley75(3666)

P(8,5)
_____________
P(3,1)x P(5,4)
8P5=6,720
3P1=3
5P4=120
6,720/3*120=6,720/360=18.6667 ANSWER.